OpenStudy (anonymous):

shortest distance between the lines (x+1)/7=(y+1)/-6=z+1/1 and (x-3)/1=(y-5)/-2=(z-7)/1

OpenStudy (anonymous):

@amistre64 plz help!

OpenStudy (amistre64):

if no one else comes along, ill be able to look at this after classes, which is in about 5 hours :/

OpenStudy (anonymous):

okay! will await ur answer!! Big thanks in advance!

OpenStudy (amistre64):

ok, im free again for the moment :)

OpenStudy (amistre64):

the shortest distance is going to be when the line (a vector) between them is perp to both; to get a perp vector we have to cross the 2 given line vectors

OpenStudy (amistre64):

(x+1)/7=(y+1)/-6=z+1/1; <7,-6,1> (x-3)/1=(y-5)/-2=(z-7)/1; <1,-2,1> x 7 1 x:-6--2 y -6 -2 -y: 7 - 1 z 1 1 z:-14--6 perp vector to both lines is: <4,6,8>, or reduced to <2,3,4> if need be

OpenStudy (amistre64):

so, what im thinking now is that we can use this perp to define the planes that the lines live in; and then we can determine the distance between the planes as the distance between the lines

OpenStudy (amistre64):

(x+1)/7 (y+1)/-6 (z+1)/1 :point is (-1,-1,-1) (x-3)/1 (y-5)/-2 (z-7)/1; point is (3,5,7) or as a side observation, it looks like we can just take those xyz parts and nestle them into the perp vector :) 2(x+1)+3(y+1)+4(z+1)=0 2x+3y+4z +10=0 2(x-3)+3(y-5)+4(z-7)=0 2x+3y+4z - 49 = 0

OpenStudy (amistre64):

now, there is a formula floating out there that you can use if you want to to find the distance between the planes, but im not much for formulas/

OpenStudy (amistre64):

one idea i have is to go about this similar to the last one i help you one, take the perp vector and a point from one of the planes, and see where it goes thru the other plane, and t will be the distance between them

OpenStudy (amistre64):

oh, and t will be the scalar of the vector, so the length of the vector 2t,3t,4t will be the distance x=-1+2t y=-1+3t z=-1+4t 2(-1+2t-3)+3(-1+3t-5)+4(-1+4t-7)=0 -8 + 4t -18 +9t-32+16t=0 t(29) = 58 t = 58/29

OpenStudy (amistre64):

distance: sqrt(4t^2+9t^2+16t^2) sqrt(t^2(4+9+16)) t sqrt(29)

OpenStudy (amistre64):

you might wanna go thru and make sure my concepts are sound, and that i didnt mismath it along the way ;)